OpenStudy (anonymous):

what is the sq root of 16W^6 Assume that the variable represents a positive real number.

OpenStudy (anonymous):

If you choose W=1 that would give you 16(1)^6 = 16(1) square root of 16 is 4. 4 x 4 = 16 Make sense?

OpenStudy (alekos):

thats correct tophat but if we leave w as it is we have the following sqrt(16w^6) = sqrt(16).sqrt(w^6) = 4w^3

OpenStudy (anonymous):

Ahh, gotcha...yeah it's hard to tell on here the level of math, sorry for dumbing it down :)